A car travels at 40 km/h for the first 30 minutes, then at 55 km/h for the next hour. What total distance (in km) did the car travel?
Answer by TimothyLamb(4379)

s = d/t
d = st
---
distance of first leg:
a = 40*(30/60)
---
distance of second leg:
b = 55*1
---
a + b = total distance
a + b = 40*(30/60) + 55*1
40*(30/60) + 55*1
---
copy and paste the above expression in to this calculator:
http://sooeet.com/math/scientific-calculator.php
---
75 km
---
Free algebra tutoring live chat:
http://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
http://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
http://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
http://sooeet.com/math/system-of-linear-equations-solver.php
---